OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. -6y = -12 -6y = 12 -2y = -12 2y = 12

OpenStudy (anonymous):

3x -4y=7 3x+2y=-5 -6y=12 The Answer is B.